A mass of 100 kg is being pushed along the floor at a constant velocity of 2 m/s. If the coefficient of kinetic friction is 0.25, at what rate is work (power) being done in watts, in horsepower?


1
Expert's answer
2021-05-26T13:02:54-0400


Gives

Mass(m)=100kg

"\mu=0.25"

Velocity (v)=2"m\/sec"

"F=\mu R=\mu mg"

F="0.25\times100\times 9.8=245N"

Power

"P=F.v"

"\frac{dW}{dt}=P=245\times2\cos0°"

"\frac{dW}{dt}=P=490W"

"\frac{dW}{dt}=P=\frac{490}{746} hp=0.6568hp"

1 hp=746W

"P=\frac{dW}{dt}=0.6568hp"